# IJCAI–25 Formatting Instructions

### v1

@@ Abstract
Conventional retrieval-augmented neural machine translation (RANMT) systems leverage bilingual corpora, e.g., translation memories (TMs). Yet, in many settings, in-domain monolingual target-side corpora are often available. This work explores ways to take advantage of such resources by retrieving relevant segments directly in the target language, based on a source-side query. For this, we design improved cross-lingual retrieval systems, trained with both sentence level and word-level matching objectives. In our experiments with two RANMT architectures, we first demonstrate the benefits of such cross-lingual objectives in a controlled setting, obtaining translation performances that surpass standard TM-based models. We then showcase our method on a real-world set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements of our new techniques, which outperform both the baseline setting, and general-purpose cross-lingual retrievers.
@@ 1 Introduction
The use of Retrieval-Augmented Generative Models (Li et al., 2022) is rapidly expanding, owing to their built-in ability to condition generations with illustrative segments retrieved from memory. The combination of (machine) retrieval and (human) regeneration has long been a key principle of Computer Aided Translation (CAT) systems (Arthern, 1978; Kay, 1997; Bowker, 2002): segments resembling the source sentence are first retrieved from Translation Memories (TM), providing translators valuable suggestions that can then be edited into a new translation. Such techniques have been transposed in Statistical MT (Koehn and Senellart, 2010), more recently in Retrieval-Augmented Neural Machine Translation (RANMT) e.g., (Gu et al., 2018), making the entire process fully automatic. Most subsequent work has continued to leverage parallel data in the form of TMs, performing retrieval in the source language, then using the retrieved target side text in generation. Similar principles underlie few-shot MT with large language models, where both the source and target sides of examples are inserted into the prompt Brown et al. (2020).
@@
TM-based approaches typically retrieve examples with lexical fuzzy matchers, using BM25 and/or the Levenshtein distance (LED) as filters or rankers (Bouthors et al., 2024). These are known to be computationally efficient, and to often surpass continuous-space retrievers (Xu et al., 2020): this is because lexical (source) matches often come with appropriate translation of terms. However, relying on source-side lexical similarity raises two issues that can cause irrelevant segments to be retrieved: (a) noisy alignments between source and target side of a parallel instance; (b) intrinsic morphosyntactic divergences between the two languages (Dorr, 1994).
@@
These problems can be sidestep by performing retrieval directly on the target side, using Cross-Lingual Information Retrieval (CLIR) techniques. An important additional benefit would be to dispense with the use of parallel data and to potentially inform the generation process with relevant monolingual resources. As demonstrated by the recurrent use of target side monolingual data thanks to back-translation Sennrich et al. (2016), such resources are much easier to find than parallel ones in many domains, and also devoid of so-called translationese phenomena Bogoychev and Sennrich (2020). As such, monolingual corpus constitute a very useful, yet under-studied, source of high-quality segments for RANMT. To search such resources, cross-lingual retrievers are readily available thanks to multilingual sentence encoders such as LASER (Artetxe and Schwenk, 2019b) or LaBSE (Feng et al., 2022). They however lack the critical ability to ensure that the retrieved examples will have a large lexical overlap with the source text.
@@
In this work, we propose novel methods to boost the lexical matching abilities of language-agnostic sentence encoders, thereby improving their effectiveness for RANMT. For this, we consider various ways to fine-tune a pre-trained encoder into reproducing the behaviour of lexical fuzzy matchers. We experiment with two language pairs (English-French and German-English) and multiple domains of various sizes and diversities. We consider two types of RANMT systems: an autoregressive source-augmented transformer (Bulte and Tezcan, 2019) and a non-autoregressive edit-based model (Bouthors et al., 2023). We first study the performance of CLIR techniques in a controlled setting, showing that they match standard TM-based models with source-side retrieval. We then consider a much larger set-up, where the target monolingual resources far exceed the amount of parallel data and observe large improvements using our new techniques, which outperform both the baseline setting, as well as generic cross-lingual retrievers.
@@ 2 Related Work
Our work falls under the scope of Retrieval-Augmented Generation, leveraging both lexical (BM25 (Robertson and Walker, 1994), Levenshtein distance (Levenshtein, 1965)) and semantic methods for the Information Retrieval (IR) component. Semantic methods usually rely on dual encoders trained on parallel data (Gillick et al., 2018) with a contrastive loss (Sohn, 2016). Our focus is on massively multilingual sentence encoders that can encode mutual translations as close neighbors, such as e.g, LASER (Artetxe and Schwenk, 2019b) and LaBSE (Feng et al., 2022).
In recent years, there has been a growing focus on integrating retrieval components into NMT systems. One key motivation is to enhance model transparency by providing users with the retrieved examples that inform the generated output (Rudin, 2019). A simple implementation of this idea encodes the target side of the example(s) along with the source sentence, providing an additional translation context (Bulte and Tezcan, 2019; Gu et al., 2018; Xia et al., 2019; He et al., 2021; Cheng et al., 2022; Agrawal et al., 2023). It is also possible to leverage both the source and target sides of the example(s) (Pham et al., 2020; Reheman et al., 2023).
Instead of regenerating a complete text with an extended context, edit-based models aim to perform minimal changes to the retrived examples and generate their output in a non-autoregressive (NAR) fashion. Such techniques are studied in (Niwa et al., 2022; Xu et al., 2023; Zheng et al., 2023), adapting the Levenshtein transformer (Gu et al., 2019). This approach is extanded to handle multiple examples that are simultaneously edited in (Bouthors et al., 2023).
MT systems based on Large Language Models (LLMs) can also seamlessly accomodate examples though In-Context Learning (ICL), where the LLM prompt is enriched with multiple demonstrations of the translation task, comprising both the source and target sides of parallel sentences (Radford et al., 2019). Several studies have focused on optimizing the performance of such architectures for NMT, examining the impact of prompt modifications, the number of demonstrations and the retrieval procedure (Moslem et al., 2023; Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023; Bawden and Yvon, 2023; Zebaze et al., 2024).
Our lexical alignment objective corresponds to a form of neural metric learning, where a high dimensional distance space is reduced to a smaller one (Suárez et al., 2021). This field of research aims to learn distances from the data (Kulis, 2013; Bellet et al., 2015), which is particulary useful for similarity-based algorithms used in IR (Cakir et al., 2019).
@@ 3.1 Problem Formulation
Given a source sentence $\mathbf{x}$ and a corpus of segments $\mathcal{D}=(\mathbf{y}_{1},\dots,\mathbf{y}_{N})$ in the target language, the goal of CLIR in NMT is to retrieve $k$ relevant examples: (~$(\tilde{}\mathbf{y}_{1},\dots,\tilde{}\mathbf{y}_{k})$.
Our approach closely follows neural-based cross-lingual retrievers, and comprises two main steps. First, $\mathcal{D}$ is encoded with a multilingual sentence encoder $E_{\theta}$. Then, for any query source sentence $\mathbf{x}$, retrieval is performed as a $k$NN similarity search in the embedding space, using the cosine similarity:
@@
The quality of the retrieved sentences only depends on the choice of $E_{\theta}$. Usually, $E_{\theta}$ is trained as a Siamese encoder with a contrastive loss on parallel data, a usual IR objective (Sohn, 2016) that primarily captures semantic similarity. However, it is desirable for the retrieved $\tilde{\mathbf{y}}_{i}$ instances to exhibit surface-level similarities to what could potentially be a translation $\mathbf{y}$ of the source $\mathbf{x}$. In specialized domains (like law or medicine), adherence to domain-specific terminology takes precedence over semantic similarity, thereby motivating the design of our proposed lexical-based alignment objective. To address this need, we study various fine-tuning objectives based on the Levenshtein similarity. These objectives are specifically designed to boost the retrieval of segments with a large numbers of lexical matches in the source query, by learning a new embedding space, achieved through the optimization of parameters $\theta$, such that:
@@
$\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})=1-\frac{\Delta(\mathbf{y},\tilde{}\mathbf{y})}{\max(|\mathbf{y}|,|\tilde{}\mathbf{y}|)},$ (3)
where $\Delta$ is the Levenshtein distance (Levenshtein, 1965).
@@ 3.2 Training
We consider three objective functions. The first two learn to predict the Levenshtein similarity. We choose $f$ to be a learnable fonction that maps $[-1,1]$ to $[0,1]$:
@@
$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y})=\operatorname{Err}(f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})),\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})),$ (5)
where the error function $\operatorname{Err}$ is either the mean square error (MSE) or the mean absolute error (MAE) defined as:
@@
The third objective is a “learning to rank” objective (Cao et al., 2007). Given a set of $k$ retrieved segments ~$\tilde{}\mathbf{y}_{[1:k]}$ sorted in decreasing order w.r.t. $\operatorname{Lev}(\mathbf{y},\cdot)$, it computes:
$\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y}_{[1:k]})=$ $\sum_{i>j}\max(0,\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{i})-\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{j})+$
$m\times|\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})-\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})|)$ (8)
with $m$ a scaling factor. Equation (8) defines a pairwise margin loss, [^1] Derived from the triplet loss (Schultz and Joachims, 2003). with an adaptive margin proportional to the absolute difference between the two individual $\operatorname{Lev}$ similarities.
@@ 4.1 Data
We consider two translation tasks: from English to French (en-fr) and from German to English (de-en). Each test corpus contains a wide variety of textual domains (16), all from the Opus corpus (Tiedemann, 2012). The details are shown in appendix A.
We also prepare a large corpus of Wikipedia pages in French. [^2] https://huggingface.co/datasets/Plim/fr_wikipedia_processed. We split the text in pseudo-sentences via regular expression matching and remove duplicates as well as fuzzy matches from the valid/test sets [^3] Each best match with $\operatorname{Lev}>0.9$ is simply removed. of the parallel Wikipedia corpus split, ending up with about $45$M French segments.
@@ 4.2 Metrics
The success of cross-lingual retrieval is assessed with the average target side Levenshtein similarity (eq. (3)) between the 1-best match and the reference, for sentences in the validation set, with retrieval performed from the entire train set. We additionally report xsim error rates, which measure the abililty to identify matched sentences in a bilingual parallel corpus (Artetxe and Schwenk, 2019a).
We assess translation quality with BLEU scores (Papineni et al., 2002) computed with SacreBLEU (Post, 2018),444signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.1.0; and COMET-20555Unbabel/wmt22-comet-da, using the official implementation. scores (Rei et al., 2020).
@@ 5.1 Retrieval Settings
Our controlled experiments use parallel corpora in two language pairs (en-fr and de-en). This allows us to perform a head-to-head comparison between standard source-side fuzzy matching and our novel cross-lingual retrieval methods. In each configuration, we retrieve up to $k=3$ closest examples. We compare the following retrieval methods:
  • fuzzy-src: source-to-source fuzzy matching ranked by their $\operatorname{Lev}$ similarity, as in standard RANMT settings;
  • fuzzy-gold: target-to-target fuzzy matching, using the same procedure as fuzzy-src. This setup requires access to reference target translations and corresponds to an oracle score, giving an empirical upper-bound of what an ideal retriever could achieve;
  • fuzzy-bt: the target side of training segments are back-translated into the source language Sennrich et al. (2016), then used for retrieval as in fuzzy-src. This simulates an alternative to cross-lingual retrieval in situations where only target side examples are available;
  • dense: a BERT base Devlin et al. (2019) model trained from scratch with a contrastive loss (Sohn, 2016), usable for source-to-target retrieval. Given a batch $\mathcal{B}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n}))$, the loss is computed as:
    $\mathcal{L}(\mathcal{B})=-\sum_{i}\log\frac{\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{i}))}{\sum_{j}\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{j}))};$ (9)
  • dense+bow: we add a bag-of-word loss to dense as in (Cai et al., 2021) to enforce lexical information in cross-lingual encodings. Given a pair of parallel sentences $(\mathbf{x},\mathbf{y})$, the loss is computed as:
    $\mathcal{L}(\mathbf{x},\mathbf{y})=-\sum_{w\in\mathcal{Y}}\log p_{\theta_{1}}(w|\mathbf{x})-\sum_{w\in\mathcal{X}}\log p_{\theta_{2}}(w|\mathbf{y}),$ (10)
    where $\mathcal{X}$, $\mathcal{Y}$ are the bag-of-words of $\mathbf{x}$ et $\mathbf{y}$; $p_{\theta_{1}}$ and $p_{\theta_{2}}$ are distinct linear projections of $E_{\theta}(\mathbf{x})$ and $E_{\theta}(\mathbf{y})$;
@@
  • LaBSE: Similar to dense, LaBSE (Feng et al., 2022) uses a dual encoder model based on the BERT base architecture and is trained with a contrastive loss; it additionnally benefits from pre-training on large sets of monolingual data, and covers more than 110 languages;
[…]
In our setting, fine-tuning updates all the model parameters. The global learning rate is 1e-4, as for $a$ and $b$ (eq. 4) it is set to 1e-2. Each source $\mathbf{x}$ is presented with three ~$\tilde{}\mathbf{y}$ (eq. 5) determined by the top 3 scoring matches obtained with dense+bow. The validation score is an in-batch normalized discounted cumulative gain (NDCG) score (Wang et al., 2013).
The retrieval is always performed in an in-domain setting. For all methods, a threshold is further applied to filter out low similarity segments. For each retrieval method, we adjust this threshold on the validation set to achieve an averaged retrieval rate [^6] Proportion of sentences with at least one retrieved example. of $50$%. Using a constant retrieval rate for all retrievers ensures that no system is unduly rewarded for retrieving a greater number of similar examples, and that our comparisons of retrievers only reflects differences in the quality of the TM matches. Please, refer to appendix B for the complete search configurations.

### v2

# Improving Retrieval-Augmented Neural Machine Translation with Monolingual Data

@@ Abstract
Conventional retrieval-augmented neural machine translation (RANMT) systems leverage bilingual corpora, e.g., translation memories (TMs). Yet, in many settings, monolingual corpora in the target language are often available. This work explores ways to take advantage of such resources by directly retrieving relevant target language segments, based on a source-side query. For this, we design improved cross-lingual retrieval systems, trained with both sentence level and word-level matching objectives. In our experiments with three RANMT architectures, we assess such cross-lingual objectives in a controlled setting, reaching performances that match those of standard TM-based models. We also showcase our method on a real-world settings, using much larger monolingual corpora, and observe strong improvements over both the baseline setting and general-purpose cross-lingual retrievers.
Improving Retrieval-Augmented Neural Machine Translation with Monolingual Data
Maxime Bouthors   Josep Crego   Dakun Zhang   François Yvon SYSTRAN by ChapsVision, 5 rue Feydeau, F-75002 Paris, France Sorbonne Université, CNRS, ISIR, F-75005 Paris, France {mbouthors, jcrego, dzhang}@chapsvision.com francois.yvon@cnrs.fr
@@ 1 Introduction
The use of Retrieval-Augmented Generative Models (Li et al., 2022) is rapidly expanding, owing to their built-in ability to condition generations with illustrative segments retrieved from memory. The combination of (machine) retrieval and (human) regeneration has long been a key principle of Computer Aided Translation (CAT) systems (Arthern, 1978; Kay, 1997; Bowker, 2002): segments resembling the source sentence are first retrieved from Translation Memories (TM), providing translators valuable suggestions that can then be edited into a new translation. Such techniques have been transposed in Statistical MT (Koehn and Senellart, 2010), more recently in Retrieval-Augmented Neural Machine Translation (RANMT) e.g., (Gu et al., 2018), making the entire process fully automatic. Most subsequent work has continued to leverage parallel data, performing retrieval in the source language, then using the retrieved target side text in generation. Similar principles underlie few-shot MT with large language models, where both the source and target sides of examples are inserted into the prompt (Brown et al., 2020).
@@
TM-based approaches typically retrieve examples with lexical fuzzy matchers, using BM25 and/or the Levenshtein distance (LED) as filters or rankers (Bouthors et al., 2024). These are known to be computationally efficient and to often surpass continuous-space retrievers (Xu et al., 2020): this is because lexical (source) matches often come with appropriate translation of terms. However, using source-side lexical similarity raises two issues that may cause inadequate segments to be retrieved: (a) noisy alignments between source and target side of a parallel instance; (b) intrinsic linguistic divergences between the two languages (Dorr, 1994).
@@
These problems can be sidestepped by performing retrieval directly on the target side, using Cross-Lingual Information Retrieval (CLIR) techniques (Cai et al., 2021; Tamura et al., 2023). An important additional benefit is to dispense with the use of parallel data and potentially inform the generation process with relevant monolingual resources. As demonstrated by the recurrent use of target side monolingual data thanks to back-translation (Sennrich et al., 2016a), such resources are much easier to find than parallel ones in many domains, and are also devoid of so-called translationese phenomena (Bogoychev and Sennrich, 2020). As such, monolingual corpora constitute a very useful, yet under-studied, source of high-quality segments for RANMT. To search such resources, cross-lingual retrievers are readily available thanks to multilingual sentence encoders such as LASER (Artetxe and Schwenk, 2019b) or LaBSE (Feng et al., 2022). As they only rely on dense representations, such retrievers however often fail to retrieve lexically relevant examples, i.e, examples that would contain translations of the exact source words.
@@
In this work, we propose novel methods to boost the lexical matching abilities of language-agnostic sentence encoders, thereby improving their effectiveness for RANMT. For this, we consider various ways to fine-tune a pre-trained encoder into reproducing the behaviour of lexical fuzzy matchers. We experiment with three language pairs (English-French, German-English and English-Ukrainian) and multiple domains of various sizes and diversities. We consider three types of RANMT systems: an autoregressive source-augmented transformer (Bulte and Tezcan, 2019), a non-autoregressive edit-based model (Bouthors et al., 2023) and a Large Language Model (LLM), with in-context ($k$-shot) learning (Brown et al., 2020). We first study the performance of CLIR techniques in a controlled setting, showing that they match standard TM-based models with source-side retrieval. We then consider much larger setups, where the target monolingual resources far exceed the amount of parallel data and observe large improvements using our new techniques, which outperform various baseline systems, as well as RANMT architectures relying on generic cross-lingual retrievers.
@@ 2 Related Work
Retrieval Augmented NMT
Our work falls under the scope of Retrieval-Augmented Generation, leveraging both lexical (BM25 (Robertson and Walker, 1994), Levenshtein distance (Levenshtein, 1965)) and semantic methods for the Information Retrieval (IR) component. Semantic methods usually rely on dual encoders trained on parallel data (Gillick et al., 2018) with a contrastive loss (Sohn, 2016). Our focus is on massively multilingual sentence encoders that encode mutual translations as close neighbors, such as LASER (Artetxe and Schwenk, 2019b) and LaBSE (Feng et al., 2022).
In recent years, retrieval-augmented MT systems have received increasing attention. One key motivation is to enhance model transparency by providing users with the retrieved examples that inform the generated output (Rudin, 2019). A simple implementation of this idea encodes the target side of the example(s) along with the source sentence, providing an additional translation context (Gu et al., 2018; Bulte and Tezcan, 2019; Xia et al., 2019; He et al., 2021; Cheng et al., 2022; Agrawal et al., 2023). It is also possible to leverage both the source and target sides of the retrieved example(s) (Pham et al., 2020; Reheman et al., 2023).
Rather than regenerating a complete translation with extended context, edit-based models perform minimal changes to the retrieved examples and generate their output in a non-autoregressive (NAR) fashion. Such models are studied in (Niwa et al., 2022; Xu et al., 2023; Zheng et al., 2023), adapting the Levenshtein Transformer (Gu et al., 2019). Bouthors et al. (2023) generalize this approach to simultaneously handle multiple examples.
MT systems based on Large Language Models (LLMs) can also seamlessly accomodate examples through In-Context Learning (ICL), where the LLM prompt is enriched with demonstrations of the translation task, comprising both the source and target sides of parallel samples (Radford et al., 2019). Several studies have tried to optimize the performance of such architectures for MT, examining the impact of prompt modifications, the number of demonstrations and the retrieval procedure (Moslem et al., 2023; Vilar et al., 2023; Zhang et al., 2023; Hendy et al., 2023; Bawden and Yvon, 2023; Zebaze et al., 2024).
Monolingual Data in NMT
Augmenting NMT models with target side monolingual data by back-translating segments is a well-established strategy for improving model training (Sennrich et al., 2016a; Edunov et al., 2018), already considered in RANMT setup (Tezcan et al., 2024). (Cai et al., 2021; Tamura et al., 2023) dispense with back-translation, performing directly the retrieval in the target language with CLIR techniques: the latter uses existing retrievers, while the former jointly trains the retriever and the MT components.
Metric Learning
Our lexical alignment objectives (§ 3.2) corresponds to a form of neural metric learning, where a high-dimensional metric space is reduced to a smaller one (Suárez et al., 2021). This line of work aims to learn distances from the data (Kulis, 2013; Bellet et al., 2015), which is particulary useful for similarity-based algorithms used in IR (Cakir et al., 2019).
@@ 3.1 Problem Formulation
Given a source sentence $\mathbf{x}$ and a corpus of target segments $\mathcal{D}=(\mathbf{y}_{1},\dots,\mathbf{y}_{N})$, the goal of CLIR is to retrieve $k$ relevant examples: $(\tilde{}\mathbf{y}_{1},\dots,\tilde{}\mathbf{y}_{k})$.
Our approach closely follows neural-based cross-lingual retrievers, and comprises two main steps. First, $\mathcal{D}$ is encoded with a multilingual sentence encoder $E_{\theta}$. Then, for any query source sentence $\mathbf{x}$, retrieval is performed as a kkNN similarity search in the embedding space, using the cosine similarity:
@@
The quality of the retrieved sentences only depends on the choice of $E_{\theta}$. Usually, $E_{\theta}$ is trained as a Siamese encoder with a contrastive loss on parallel data, a usual IR objective (Sohn, 2016) that primarily captures semantic similarity. However, it is desirable for the retrieved $\tilde{\mathbf{y}}_{i}$ instances to exhibit lexical similarities with the unknown translation $\mathbf{y}$ of the source $\mathbf{x}$. In specialized domains (like law or medicine), adherence to domain-specific terminology takes precedence over semantic similarity, thereby motivating the design of our proposed lexical-based alignment objective. To address this need, we study various fine-tuning objectives based on the Levenshtein similarity. These objectives are specifically designed to boost the retrieval of segments with a large numbers of lexical matches in the source query, by learning a new embedding space, achieved through the optimization of parameters $\theta$, such that: [^1] As edit distances are only computed monolingually, our method does not assume any lexical overlap between $\mathbf{x}$ and $\tilde{}\mathbf{y}$.
@@
$\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})=1-\frac{\Delta(\mathbf{y},\tilde{}\mathbf{y})}{\max(|\mathbf{y}|,|\tilde{}\mathbf{y}|)}.$ (3)
$\Delta$ is the Levenshtein distance (Levenshtein, 1965).
@@ 3.2 Training
We consider three objective functions. The first two learn to predict the Levenshtein similarity (3). We choose $f$ to be a learnable function that maps [1,1][-1,1] to $[0,1]$:
@@
$\hskip-5.0pt\mathcal{L}(\mathbf{x},\mathbf{y},\tilde{}\mathbf{y})=\operatorname{Err}(f(\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y})),\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y})),$ (5)
where the error function $\operatorname{Err}$ is either the mean square (MSE) or the mean absolute error (MAE):
@@
The third objective is a “learning to rank" objective (Cao et al., 2007). Given a set of $k$ retrieved segments $\tilde{}\mathbf{y}_{[1:k]}$ sorted in decreasing order w.r.t. $\operatorname{Lev}(\mathbf{y},\cdot)$, it computes:
$\mathcal{L}($ $\mathbf{x},\mathbf{y},\tilde{}\mathbf{y}_{[1:k]})=\sum_{i>j}\max(0,U_{ij})$ (8)
$U_{ij}$ $=\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{i})-\mathrm{sim}_{\theta}(\mathbf{x},\tilde{}\mathbf{y}_{j})$
$+m\times|\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})-\operatorname{Lev}(\mathbf{y},\tilde{}\mathbf{y}_{j})|)$ (9)
with $m$ a scaling factor. Equation (9) defines a pairwise margin loss, [^2] Inspired by the triplet loss (Schultz and Joachims, 2003). with an adaptive margin proportional to the absolute difference between the two individual $\operatorname{Lev}$ similarities.
@@ 4.1 Data
We consider three translation directions: English into French (en-fr), German into English (de-en) and English into Ukrainian (en-uk). For our controlled experiments, we consider a wide variety of textual domains (16), using public parallel datasets from the Opus corpus (Tiedemann, 2012). Details are in appendix A.
For the large scale experiments, we expand the English-French setting with a large corpus of Wikipedia pages in French. [^3] https://huggingface.co/datasets/Plim/fr_wikipedia_processed. We split the text in pseudo-sentences via regular expression matching and remove duplicates as well as fuzzy matches from the valid/test sets [^4] Each best match with $\operatorname{Lev}>0.9$ is simply removed. of the parallel Wikipedia corpus, ending up with about 4545M French segments. The corresponding English-Ukrainian experiments, reported in Appendix H, reproduce the setting of Tezcan et al. (2024), using texts in the legal domain.
@@ 4.2 Metrics
The success of cross-lingual retrieval is assessed with the average target side Levenshtein similarity (eq. (3)) between the 1-best match and the reference, for sentences in the validation set, with retrieval performed from the entire train set. We also report xsim error rates, which measure the abililty to identify matched sentences in a bilingual parallel corpus (Artetxe and Schwenk, 2019a).
We assess translation quality with BLEU (Papineni et al., 2002) computed with SacreBLEU (Post, 2018),555signature: nrefs:1|case:mixed|eff:no|tok:13a| smooth:exp|version:2.1.0; as well as COMET-20666Unbabel/wmt20-comet-da, using the defaults settings. (Rei et al., 2020).
@@ 5.1 Retrieval Settings
Our experiments consider the following retrieval methods, with up to $k=3$ closest examples:
  • fuzzy-src: source-to-source fuzzy matching ranked by their $\operatorname{Lev}$ similarity, as in baseline RANMT settings;
  • fuzzy-gold: target-to-target fuzzy matching, using the same procedure as fuzzy-src. This retriever requires access to reference target translations and corresponds to an oracle setting, giving an empirical upper-bound of what an ideal system could achieve. It thus is only reported for the controlled experiments of § 6.1 and § 6.2;
  • fuzzy-bt: the target side of training segments are back-translated using NLLB (Costa-jussà et al., 2024) [^7] We used NLLB-200-distilled-1.3B out-of-the-box.
    https://huggingface.co/facebook/nllb-200-1.3B
    into the source language Sennrich et al. (2016a), then used for retrieval as in fuzzy-src. This simulates an alternative to cross-lingual retrieval in situations where only target side examples are available; [^8] Back-translation is further discussed in § G.
  • dense: a BERT base Devlin et al. (2019) model for source-to-target retrieval. It is trained from scratch with a contrastive loss (Sohn, 2016). The loss $\mathcal{L}(\mathcal{B})$ for batch $\mathcal{B}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n}))$ is defined as:
    $\sum_{i}\log\frac{\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{i}))}{\sum_{j}\exp(\mathrm{sim}(\mathbf{x}_{i},\mathbf{y}_{j}))};$ (10)
  • dense+bow: we add a bag-of-word loss to dense as in (Cai et al., 2021) to enforce lexical information in cross-lingual encodings. Given a pair of parallel sentences $(\mathbf{x},\mathbf{y})$, the loss $\mathcal{L}(\mathbf{x},\mathbf{y})$ is computed as:
    $\sum_{w\in\mathcal{Y}}\log p_{\theta_{1}}(w|\mathbf{x})-\sum_{w\in\mathcal{X}}\log p_{\theta_{2}}(w|\mathbf{y}),$ (11)
    where $\mathcal{X}$, $\mathcal{Y}$ are the bag-of-words for $\mathbf{x}$ et $\mathbf{y}$; $p_{\theta_{1}}$ and $p_{\theta_{2}}$ are distinct linear projections of $E_{\theta}(\mathbf{x})$ and $E_{\theta}(\mathbf{y})$;
@@
  • LaBSE: Similar to dense, LaBSE (Feng et al., 2022) uses a dual encoder model based on the BERT base architecture and is trained with a contrastive loss; it additionally benefits from pre-training on large sets of monolingual data, and covers more than 110 languages;
  • ft-[model]-[Err]: we fine-tune either one of the aforementioned encoders with the losses of §3.2; Err is one of $\{\text{MSE},\text{MAE},\text{Rank}\}$ (e.g. ft-LaBSE-MSE).
In our setting, fine-tuning updates all the parameters. The global learning rate is 1e-4; as for $a$ and $b$ (eq. 4) it is set to 1e-2. Each source $\mathbf{x}$ is presented with three $\tilde{}\mathbf{y}$ (eq. 5) determined by the top 3 scoring matches obtained with dense+bow. The validation score is an in-batch normalized discounted cumulative gain (NDCG) score (Wang et al., 2013).
Retrieval is always performed in an in-domain setting, ensuring that translations for a specific domain are exclusively based on examples from that domain. For all methods, a threshold is further applied to filter out low similarity segments. For each retrieval method, we adjust this threshold on the validation set to achieve an averaged retrieval rate [^9] Fraction of sentences with at least one retrieved example. of 5050%. Using a constant retrieval rate for all retrievers ensures that no system is unduly rewarded for retrieving a greater number of similar examples, so that our comparisons of retrievers only reflect differences in the quality of the TM matches. Refer to appendix B for a discussion of search configurations.